When an evacuated container whose volume is 600 mL is filled with O2 gas at 40°C and 900 mmHg, the mass of the container will increase by …………………g 


1
Expert's answer
2021-02-02T09:30:15-0500

T = 40 +273 = 313 K

V = 600 mL = 0.6 L

p = 900 mmHg = 1.18 atm

Ideal Gas Law

pV = nRT

R = 0.08206 L×atm/mol×K

"n = \frac{pV}{RT} \\

= \frac{1.18 \times 0.6}{0.08206 \times 313} \\

= 0.0275 \;mol \\

M(O_2) = 32 g/mol \\

m(O_2) = n \times M \\

= 0.0275 \times 32 \\

= 0.88 \;g"

The mass of the container will increase by 0.88 g.